OpenStudy (anonymous):

A die is rolled three times. Find the chance that the face with 4 spots appears at least one.

OpenStudy (anonymous):

Let A=probability getting 4=1/6 B=probability not getting4=1-1/6=5/6 In all three rolls of die P(not getting 4)=B*B*B=5/6*5/6*5/6=125/216 p atleast one 4=1-125/216=(216-125)/216=91/216